An object of mass m = 100g falls freely during 3sec:

a) Find the received impulse,

b) Deduce the change of the speed
1
Expert's answer
2019-06-20T11:18:25-0400

a) Impulse is the change of momentum "p=mv" of an object when the object is acted upon by a force "F"  for an interval of time "\Delta t" . The formula for impulse is:


"Impulse=Force\times time=F\cdot \Delta t"


Object falls under the force of gravity, which is


"F=mg"

where m = 100g = 0.1kg,   g=9.8m/s2 is the acceleration of gravity. Then we get


"Impulse=0.1\,kg\cdot 9.8\,m/{{s}^{2}}\cdot 3\,s=\text{2}\text{.94}\,kg\cdot \,m/s"


b) Because impulse is a measure of how much the momentum changes as a result of force acting on it for a period of time, an alternative formula for impulse is:


"Impulse=\Delta p={{p}_{final}}-{{p}_{initial}}""=m{{v}_{final}}-m{{v}_{initial}}=m\Delta v"

Then change of the speed is


"\Delta v=\frac{Impulse}{m}=\frac{\text{2}\text{,94}\,kg\cdot \,m/s}{0.1\,kg}=29.4\,m/s"